# Radar-Guided Polynomial Fitting for Metric Depth Estimation

### v2

@@ Abstract
We propose POLAR, a novel radar-guided depth estimation method that introduces polynomial fitting to efficiently transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps. Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a fundamental insight: although MDE models often infer reasonable local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale and shift (affine) transformation insufficient given three or more of these regions. To address this limitation, we use polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges. In this way, POLAR generalizes beyond affine transformations and is able to correct such misalignments by introducing inflection points. Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces local monotonicity via first-derivative regularization. POLAR achieves state-of-the-art performance across three datasets, outperforming existing methods by an average of 24.9% in MAE and 33.2% in RMSE, while also achieving state-of-the-art efficiency in terms of latency and computational cost.
@@ 1 Introduction
Metric 3D reconstruction is critical for spatial tasks such as self-driving (Maier et al., 2012; Gupta et al., 2021), where it is necessary for one to perceive the structure of the 3D environment in order to navigate. In many such systems, multiple sensors—including cameras, lidar, and radar—provide complementary information about the 3D scene. While lidar sensors offer dense and precise point clouds, they are expensive and not widely available (Raj et al., 2020). In contrast, radar sensors, particularly millimeter wave (mmWave) radars (Iizuka et al., 2003), return only about a hundred points per frame and are noisier (Han et al., 2023). Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles (Eichelberger & McCartt, 2016).
Unless camera baselines are known, images offer scaleless reconstruction. However, they are high-dimensional and sensitive to variations in illumination, object appearance, orientation, and camera viewpoint. Training robust models that can generalize across these factors often demands large-scale datasets, which are costly to collect. As a result, leveraging pretrained monocular depth estimation (MDE) foundation models (Ranftl et al., 2021; Yang et al., 2024; Bochkovskii et al., 2024; Piccinelli et al., 2025) emerges as a practical alternative. However, as monocular 3D reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or depth that lacks the fidelity needed for applications demanding accurate metric-scale reconstruction, such as mapping and navigation.
Existing approaches (Yin et al., 2023; Viola et al., 2024; Hu et al., 2024; Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025) attempt to transform these MDE predictions into accurate metric depth maps using a global affine (scale-and-shift) transformation. This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene. While effective in estimating ordinal relationships within local regions, where MDE predictions typically exhibit reasonable local (object-level) reconstructions, such linear corrections fail when multiple objects are placed at incorrect depths relative to each other. Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (Fig. 1).
@@
In this work, we challenge the common assumption that scale ambiguity in MDE is only up to an unknown global scale and shift. Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, enabling more flexible non-uniform adjustments. Our approach, POLAR, transforms the predictions of pretrained MDE models by (i) learning prototypical patterns in the configuration of radar points, (ii) establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and (iii) predicting polynomial coefficients that adaptively fit scaleless depth into metric depth. This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and cross-region misalignments. While one may adjust every prediction (pixel) of an MDE model (Li et al., 2024a), the number of degrees of freedom is the total number of pixels. Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem. On the other hand, the number of degrees of freedom in polynomial expansion is limited to the number of polynomial terms, which we empirically found to be close to the cardinality of radar point clouds, allowing our solution to be better posed and more regular. To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models.
@@ 2 Related Work
Monocular Depth Estimation. With recent advances in the scalability of neural networks (Dosovitskiy et al., 2021; Caron et al., 2021), monocular depth estimation (MDE) models can infer scaleless relative depth from a single image across diverse, unseen domains (Ranftl et al., 2020; Yang et al., 2024; Piccinelli et al., 2025; Bochkovskii et al., 2024), benefiting from large-scale datasets. Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation (Hoyer et al., 2023; Bian et al., 2019), uncertainty estimation (Poggi et al., 2020), optical flow (Zhao et al., 2020), and visual odometry (Song et al., 2023), to improve relative depth estimation. However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating absolute depth from a single image is an ill-posed problem. Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity (Viola et al., 2024). One way to address this limitation is to incorporate range-sensing modalities such as lidar (Jaritz et al., 2018; Ezhov et al., 2024), radar (Singh et al., 2023), or visual-inertial odometry (Wong et al., 2020).
@@
Image-Guided Depth Completion. Most often studied in the context of lidar-camera depth estimation, image-guided depth completion leverages the strengths of each modality: images offer dense visual context and structural priors, while lidar points provide absolute metric scale to resolve depth ambiguity. Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers (Ma & Karaman, 2018; Ma et al., 2018), late fusion, where inputs are processed by independent branches (Yan et al., 2021; Rim et al., 2025), and multi-scale fusion, which captures both local details and global scene structure (Li et al., 2020). U-Net-like architectures have been widely used for coarse-to-fine depth completion (Hu et al., 2021; Lin et al., 2022), with improvements from deformable convolutions (Park et al., 2020; Xu et al., 2020), and attention mechanisms (Rho et al., 2022; Zhang et al., 2023).
@@
Radar-Camera Depth Estimation. While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions (Raj et al., 2020). In contrast, radar provides a cost-effective alternative (Hunt et al., 2024), offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle (Paek et al., 2022; Srivastav & Mandal, 2023). The ubiquity of radar sensors (Eichelberger & McCartt, 2016) in existing automotive and robotic platforms further supports their integration into depth estimation pipelines. Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
@@
Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity (Singh et al., 2023) of radar point clouds. (Lin et al., 2020) uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map. (Long et al., 2021) leverages Doppler velocity and optical flow to associate radar points with image pixels. (Lo & Vandewalle, 2021) refines radar depth using height-extension of radar points to address elevation ambiguity. (Singh et al., 2023) introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion. (Li et al., 2024b) mitigates distribution artifacts using sparse supervision, while (Sun et al., 2024) employs a two-stage confidence-driven approach. GET-UP (Sun et al., 2025) uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features. Like our method, RadarCam-Depth (Li et al., 2024a) and TacoDepth (Wang et al., 2025) also begin with an initial MDE prediction, but refine it and decode dense depth directly from fused features, rather than employing a learned scene-fitting approach.
@@
Existing methods operate within the paradigm of completion or direct decoding of fused features, and they often rely on multi-stage training and explicit radar-pixel association learning, increasing model complexity and computational overhead. In contrast, POLAR employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale. It bypasses the need for multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
@@ 3.1 Motivation
Existing methods (Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025; Viola et al., 2024; Hu et al., 2024; Yin et al., 2023) refine MDE predictions by applying simple scale-and-shift transformations. However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths. A practical illustration of this shortcoming arises when multiple objects or local regions appear in the same scene: MDE may accurately infer the reconstruction within each object region, yet place them at incorrect relative depths with respect to each other. Such cross-region misalignments cannot be corrected by a single global scaling factor (see Fig. 1). This violates the assumption that the MDE reconstruction is up to an unknown global scale and shift.
@@ 3.2 Representation Learning and Fusion
Radar Processing. We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where each of the $N_{C}$ points is represented by its $(x,y,z)$ coordinates in three-dimensional camera space. Radar point clouds, while noisy and sparse, provide metric depth measurements. We concatenate a sinusoidal 3D positional embedding $\phi_{3D}(x,y,z)$ to $C$, and feed the resulting representation into a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
@@
Radar Aggregation. To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these $N_{P}$ prototypes capture diverse spatial and geometric properties present in the radar point cloud. Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes. Unlike existing works (Singh et al., 2023; Li et al., 2024a) that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud. This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
@@
Next, learnable prototypes are matched to the most relevant radar features for a given input. We treat the prototypes as centroids and perform soft clustering over the radar features. Each radar point is softly assigned to prototypes according to feature similarity, and the corresponding values are aggregated to yield a global scene-descriptive representation $F_{R}$:
$D_{ij}=\|P_{j}-\Phi_{r}(F_{r})_{i}\|^{2},\quad F_{R}=\text{softmax}\!\left(-D\,/\,\tau\right)\,\Psi_{r}(F_{r}),$ (2)
@@
where $\Phi_{r}$, $\Psi_{r}$ are MLPs that project the radar features $F_{r}$, the matrix $D\in\mathbb{R}^{N_{P}\times N_{C}}$ contains the pairwise squared distances between prototypes $\{P_{j}\}_{j=1}^{N_{P}}$ and projected radar features $\{\Phi_{r}(F_{r})_{i}\}_{i=1}^{N_{C}}$, and $\tau$ denotes temperature. This clustering-based formulation enables prototypes to capture recurring spatial and geometric patterns in radar point clouds. Comparatively, applying lidar depth completion methods, which densify a sparse projection of points using surrounding context, to a radar point cloud results in poor performance (see Sec. E). This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution (Singh et al., 2023).
where $E\in\mathbb{R}^{(H\times W)\times c_{z}}$ is a learned 2D positional embedding, and $\Phi_{R}$, $\Psi_{R}$ are MLPs that project the aggregated radar features $F_{R}$. $c_{z}$ is set equal to $c_{r}$ to align the depth and radar features within a common embedding space, facilitating cross-modal fusion, and $S$ is reshaped to be in $\mathbb{R}^{H\times W\times c_{z}}$.
$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (4)
3.3 Polynomial Fitting
This is in contrast to directly mapping to a metric-scale depth map (Singh et al., 2023; Li et al., 2024a), where the degrees of freedom comprise every predicted pixel. Each corresponding point and pixel prediction can be viewed as a constraint or equation in our estimation system. As the radar point cloud ($\sim 10^{2}$ points) only occupies a small subset of the image space ($\sim 10^{6}$ pixels), the solution is underdetermined and often irregular (see Fig. 4). Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab. 6). POLAR offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
In terms of overhead, compared to predicting two coefficients (scale and shift) for an affine transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost ($0.01\%$ more FLOPs): the final MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Sec. B for the complete derivation).
By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates $z$. Lower-order coefficients ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) establish a broad global scale, providing a metric basis for the initially scaleless $z$. Higher-order coefficients ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) introduce both low-frequency corrections, e.g., resolving cross-object misalignments (Fig. 1), as well as high-frequency corrections, e.g., sharpening fine object boundaries (Fig. 4, Image C).
Table 1: Quantitative results on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft test sets evaluated with various maximum evaluation distances.

### v3

@@ Abstract
We propose POLAR, a novel radar-guided depth estimation method that introduces polynomial fitting to efficiently transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps. Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a fundamental insight: although MDE models often infer reasonable local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale and shift (affine) transformation insufficient given three or more of these regions. To address this limitation, we use polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust predictions non-uniformly across depth ranges. In this way, POLAR generalizes beyond affine transformations and is able to correct such misalignments by introducing inflection points. Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces local monotonicity via first-derivative regularization. POLAR achieves state-of-the-art performance across three datasets, outperforming existing methods by an average of 24.9% in MAE and 33.2% in RMSE, while also achieving state-of-the-art efficiency in terms of latency and computational cost.
@@ 1 Introduction
Metric 3D reconstruction is critical for spatial tasks such as self-driving [maier2012real, GUPTA2021100057], where it is necessary for one to perceive the structure of the 3D environment in order to navigate. In many such systems, multiple sensors—including cameras, lidar, and radar—provide complementary information about the 3D scene. While lidar sensors offer dense and precise point clouds, they are expensive and not widely available [raj2020survey]. In contrast, radar sensors, particularly millimeter wave (mmWave) radars [iizuka2003millimeter], return only about a hundred points per frame and are noisier [han20234d]. Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles [eichelberger2016toyota].
@@
Unless camera baselines are known, images offer scaleless reconstruction. However, they are high-dimensional and sensitive to variations in illumination, object appearance, orientation, and camera viewpoint. Training robust models that can generalize across these factors often demands large-scale datasets, which are costly to collect. As a result, leveraging pretrained monocular depth estimation (MDE) foundation models [ranftl2021vision, depth_anything_v2, depth-pro, piccinelli2025unidepthv2] emerges as a practical alternative. However, as 3D reconstruction from a single 2D image is inherently ill-posed, they typically infer scaleless relative depth, or depth that lacks the fidelity needed for applications demanding accurate metric-scale reconstruction, such as mapping and navigation.
Existing approaches [yin2023metric3d, viola2024marigold, hu2024depthcrafter, zeng2025rsa, ding2025fixing, yu2025relative] attempt to transform these MDE predictions into accurate metric depth maps using a global affine (scale-and-shift) transformation. This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene. While effective in estimating ordinal relationships within local regions, where MDE predictions typically exhibit reasonable local (object-level) reconstructions, such linear corrections fail when multiple objects are placed at incorrect depths relative to each other. Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (Fig. 1).
In this work, we challenge the common assumption that scale ambiguity in MDE is only up to an unknown global scale and shift. Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, enabling more flexible non-uniform adjustments. Our approach, POLAR, transforms the predictions of pretrained MDE models by (i) learning prototypical patterns in the configuration of radar points, (ii) establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and (iii) predicting polynomial coefficients that adaptively fit scaleless depth into metric depth. This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and cross-region misalignments. While one may adjust every prediction (pixel) of an MDE model [li2024radarcam], the number of degrees of freedom is the total number of pixels. Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem. On the other hand, the number of degrees of freedom in polynomial expansion is limited to the number of polynomial terms, which we empirically found to be close to the cardinality of radar point clouds, allowing our solution to be better posed and more regular. To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models.
@@ 2 Related Work
Monocular Depth Estimation. With recent advances in the scalability of neural networks [dosovitskiy2020image, caron2021emerging], monocular depth estimation (MDE) models can infer scaleless relative depth from a single image across diverse, unseen domains [ranftl2020towards, depth_anything_v2, piccinelli2025unidepthv2, depth-pro], benefiting from large-scale datasets. Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation [hoyer2023improving, bian2019unsupervised], uncertainty estimation [poggi2020uncertainty], optical flow [zhao2020towards], and visual odometry [song2023unsupervised], to improve relative depth estimation. However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating depth from a single image is an ill-posed problem. Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity [viola2024marigold]. One way to address this limitation is to incorporate range-sensing modalities such as lidar [jaritz2018sparse, ezhov2024all], radar [singh2023depth], or visual-inertial odometry [wong2020unsupervised].
@@
Image-Guided Depth Completion. Most often studied in the context of lidar-camera depth estimation, image-guided depth completion leverages the strengths of each modality: images offer dense visual context and structural priors, while lidar points provide absolute metric scale to resolve depth ambiguity. Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers [sparse_to_dense, Ma2018SelfSupervisedSS], late fusion, where inputs are processed by independent branches [RigNet, rim2025protodepth], and multi-scale fusion, which captures both local details and global scene structure [msg_chn]. U-Net-like architectures have been widely used for coarse-to-fine depth completion [Hu2021PENetTP, Lin2022DynamicSP, chen2024uncle], with improvements from deformable convolutions [Park2020NonLocalSP, Xu2020DeformableSP], and attention mechanisms [Rho2022GuideFormerTF, CompletionFormer].
@@
Radar-Camera Depth Estimation. While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions [raj2020survey]. In contrast, radar provides a cost-effective alternative [hunt2024radcloud], offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle [paek2022k, srivastav2023radars]. The ubiquity of radar sensors [eichelberger2016toyota] in existing automotive and robotic platforms further supports their integration into depth estimation pipelines. Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
@@
Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity [singh2023depth] of radar point clouds. [lin2020depth] uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map. [long2021full] leverages Doppler velocity and optical flow to associate radar points with image pixels. [lo2021depth] refines radar depth using height-extension of radar points to address elevation ambiguity. [singh2023depth] introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion. [li2023sparse] mitigates distribution artifacts using sparse supervision, while [sun2024cafnet] employs a two-stage confidence-driven approach. GET-UP [sun2024get] uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features. Like our method, RadarCam-Depth [li2024radarcam] and TacoDepth [wang2025tacodepth] also begin with an initial MDE prediction, but refine it and decode dense depth directly from fused features, rather than employing a learned scene-fitting approach.
@@
Existing methods operate within the paradigm of completion or direct decoding of fused features, and they often rely on multi-stage training and explicit radar-pixel association learning, increasing model complexity and computational overhead. In contrast, POLAR employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale. It bypasses the need for multi-stage systems that learn explicit correspondences while enabling flexible depth corrections.
@@ 3.1 Motivation
Existing methods [zeng2025rsa, ding2025fixing, yu2025relative, viola2024marigold, hu2024depthcrafter, yin2023metric3d] refine MDE predictions by applying simple scale-and-shift transformations. However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths. A practical illustration of this shortcoming arises when multiple objects or local regions appear in the same scene: MDE may accurately infer the reconstruction within each object region, yet place them at incorrect relative depths with respect to each other. Such cross-region misalignments cannot be corrected by a single global scaling factor (see Fig. 1). This violates the assumption that MDE reconstruction is up to an unknown global scale and shift.
@@ 3.2 Representation Learning and Fusion
Radar Processing. We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where each of the $N_{C}$ points is represented by its $(x,y,z)$ coordinates in three-dimensional camera space. Radar point clouds, while noisy and sparse, provide metric depth measurements. We concatenate a sinusoidal 3D positional embedding $\phi_{3D}(x,y,z)$ to $C$, and feed it into a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
@@
Radar Aggregation. To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these $N_{P}$ prototypes capture diverse spatial and geometric properties present in the radar point cloud. Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes. Unlike existing works [singh2023depth, li2024radarcam] that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud. This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
@@
Next, learnable prototypes are matched to the most relevant radar features. We treat the prototypes as centroids and perform soft clustering over the radar features. Each radar point is softly assigned to prototypes according to feature similarity, and the corresponding values are aggregated to yield a global scene-descriptive representation $F_{R}$:
$D_{ij}=\|P_{j}-\Phi_{r}(F_{r})_{i}\|^{2},\quad F_{R}=\sigma\!\left(-D\,/\,\tau\right)\,\Psi_{r}(F_{r}),$ (2)
@@
where $\Phi_{r}$, $\Psi_{r}$ are MLPs that project the radar features $F_{r}$, the matrix $D\in\mathbb{R}^{N_{P}\times N_{C}}$ contains the pairwise squared distances between prototypes $\{P_{j}\}_{j=1}^{N_{P}}$ and projected radar features $\{\Phi_{r}(F_{r})_{i}\}_{i=1}^{N_{C}}$, and $\sigma$ denotes softmax. This clustering-based formulation enables prototypes to capture recurring spatial and geometric patterns in radar point clouds. Comparatively, applying lidar depth completion methods, which densify a sparse projection of points using surrounding context, to a radar point cloud results in poor performance (see Supp. Mat.). This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution [singh2023depth].
@@
Table 1: Quantitative results on nuScenes, ZJU-4DRadarCam, and View-of-Delft test sets evaluated with various maximum depths.
where $E\in\mathbb{R}^{(H\times W)\times c_{z}}$ is a learned 2D positional embedding, and $\Phi_{R}$, $\Psi_{R}$ are MLPs that project the aggregated radar features $F_{R}$. $c_{z}$ is set equal to $c_{r}$ to align the depth and radar features within a common embedding space, facilitating cross-modal fusion, and $S$ is reshaped to $H\times W\times c_{z}$.
$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (4)
Figure 3: Qualitative results on nuScenes. GET-UP and RadarCam-Depth (RC-D) fail to reconstruct entire regions, yielding objects with large depth errors. Raw MDE yields reasonable relative reconstructions but suffers from incorrect global scale and cross-object misalignments. POLAR leverages polynomial fitting to recover a global scale and correct these misalignments.
3.3 Polynomial Fitting
This contrasts with directly mapping to a metric-scale depth map [singh2023depth, li2024radarcam], where the degrees of freedom comprise every predicted pixel. Each prediction of corresponding points and pixels can be viewed as a constraint or equation in our estimation system. As the radar point cloud ($\sim 10^{2}$ points) only occupies a small subset of the image space ($\sim 10^{6}$ pixels), the solution is underdetermined and often irregular (see Fig. 3). Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab. 4). POLAR offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
In terms of overhead, compared to predicting the scale and shift for an affine transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost ($0.01\%$ more FLOPs): the final MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Supp. Mat. for the complete derivation).
By introducing polynomial terms, our method adjusts depth according to the initial depth estimates $z$. Lower-order coefficients ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) establish a broad global scale, providing a metric basis for the initially scaleless $z$. Higher-order coefficients ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) introduce both low-frequency corrections, e.g., resolving cross-object misalignments (Fig. 1), as well as high-frequency corrections, e.g., sharpening fine object boundaries (Fig. 3, Image C).